Agenda/Minutes on 2021/1/29
CRY+VIS Interface Meeting
Participants: Takahashi, Kimura, Ikeda, Tomaru, Ushiba, Miyo, Hirata, Aso, Sato
- Schedule for IX: Connection is scheduled in the week of 2/8.
- Schedule for IY: Disconnection was scheduled on 3/11.
- Schedule for EX: Connection was scheduled in the week of 3/15.
- Corrugate tube fixing jig: Re-measurement is planed in 2/26-3/1 at IYV.
- Commissioning in vacuum: Two phased commissioning will be tried in IX at beginning. After that, one phase is default.
- Acceptance check: Cryo-payload in IX and Type-A tower in EY are on going.
- Jig for FC painting: Under manufacturing.
- PF actuator: by the end of January.
[Miyo] Although this project must start after we define the actuator efficiency of PF, I (Miyo) have not defined yet. Progress is here.
Participants: Takahashi, Fabian, Hirata, Miyo, Sato, Sugiyama, Washimi, Terrence, Aso, Tanaka, Ishizaki
[Miyo] All VIS works in this week are listed here.
In order to resolve unstable network trouble of picomotor drivers that have been going on for a long time, Ikeda-san is investigating the network connection under the DGS network. He have completed to make a network diagram in this week. Report for the investigation is uploaded in JGW-E2112439. Because we found the network is too messy to manage, we separate the network for the drivers including stepper motors from the other equipments e.g. workstation. So, we are planning to make new independent network for the drivers. In this new network, drivers and script machine named k1script (now we are change to k1script1 which can run Python3 environment) will be only connected. We now think we can improve the network trouble should be resolved.
Ikeda-san found that the toggle switch of the stepper motor driver was turned on, which means that the motor driver works as "test mode". According to his statement, in this mode, we can not turn off the driver power in remotely. In fact, if the switch is turned up meaning the test mode, BIO signal is shorted to the ground according to JGW-D1605366-v1. Then, he checked the status of the switch in all area. He found that some switch was on and turn off these switch for remote operation klog#15854. Additionally, BS wasn't even cabling according to Ikeda-san. It should be done.
[Fabian] We inspected SR3 payload and chamber and came up with a plan for installing a thermistor.
- There's a mechanical design already.
- We have the thermistors already.
- The plan is to install them in 7 suspensions.
- Type-Bp chambers have unused connectors we can use for this application.
- Show diagram of how this would look like in Type-B. The diagram for Type-Bp still has to be done.
- Implemented the cabling for cryo-payload in IX.
- Assembled the connector cables for ACCs.
- Assembled the geophone pods for ITMY.
We checked the traverser cables. See klog report 15836
- We identified cable R2 as the one failing. We used a multimeter to check for electrical continuity. In R1, T1 and T2 cables there was continuity between pins 1 and 9, and between pins 2 and 3. However, in R2 there was only continuity between pins 2 and 3.
- We removed the in-vacuum cable and checked it with the multimete, but it was fine. We checked the motor also.
We realized then it was the in-air connector adapter the one failing. We borrowed one from another place and it worked well. See in this picture a thread is broken in the adapter.
- We should make more adapters and make sure the traverser works well.
Gluing of the wire. See klog 15851.
- We set up an oplev using a laser level.
- We locked the payload. In the case of the vertical bars locking the RM to the IM, we used only one out of two becuase a cable was in the way of a locking screw.
- The screws in the front ring came out smothly.
We applied adhesive as shown in this diagram. See pictures in the album.
- However, it's been two days later and the adhesive hasn't cured asin previous times.
- I checked the pictures and I used the correct proportions. We used an improvised UHV compatible Al tray.
- We mixed again in order to test the adhesive and we will check on Monday:
- We used a plastic Petri dish and a small baking Al tray provided by Tomasru-san.
- The new mixtures look translucent, as expected from previous experience, but the one we used on Wednesday looks more transparent.
- Maybe some of the component with low viscosity was trapped in the folds of the tray.
- How should we proceed? Let us discuss.
- Yet to do once we finish
- Shake the suspension to check if the mirror changes pitch as before?
- Measure the orientation of the mirror with respect to the horizontal with an oplev, using the IM picos set the pitch to 1/300 and adjust the OSEM if necessary.
- Move the traverser while inspecting visually.
Stepper motor for SF GAS might be stucked klog#15857. So we should open the top chamber.
The SC document is circulate in the previous chief meeting.
I will start the suspension commissioning for Type-Bp after the installation of the OpLev.
- I will make DC control for SR3 because Ushiba-san needs to align the green beam to ITMY on next Monday. I will make the control in weekend.
- I'm continuing to make a simulation model by studying the many unorganized files left by people in the past..
- Measured the TFs in both original LVDT driver and new one. 50dB improvement was found at 100Hz.
- The improvement will be confirmed using the BF damper in PRM.
- Optimal capacitance would make the trans power increase.
- Type-B/Bp remedy work is on going.
- Installation of Type-A is on going in EY.
- Procedure documents.
- Criteria in the acceptance test.
- O3GK Noise Budget paper [Washimi] : The paper writing team (Kokeyama, KYamamoto, Washimi) is writing the draft.
- Seismic noise through Type-A tower : on going
This noise is evaluated by the TF (Shoda-san's simulation). -> Does it include the actual situation of GAS filters? -> Maybe "yes" (by Miyo-kun)
- Type-A ctrl : on going
- Type-B ctrl : almost written. Please check it.
- Type-Bp ctrl : almost written. Please check it.
- Seismic noise through Type-A tower : on going
- Temperature setting around Type-B. (please include all suspensions not only Type-B but also Type-A, Bp)
GAS drift include O3GK (klog)
The week from 2/1
- Takahashi: 2/1-5
- Sato: 2/1-5
- Hirata: 2/1-5
- Yano: 2/1-5
The week from 2/8
- Takahashi: 2/8-12
- Sato: 2/8-12
- Hirata: 2/8-12
- Yano: 2/8-12
- On 2020/2/5(Fri)